OpenStudy (anonymous):

1/x^2 + 1/x =6 solve for x

OpenStudy (anonymous):

make the left side into 1 equation.

OpenStudy (asnaseer):

multiply both sides by\(x^2\) and solve the resulting quadratic equation.

OpenStudy (anonymous):

\[x^2/ x^2 + x^2/x^2 = 6x^2\]

OpenStudy (anonymous):

2x^2 = 6x^2

OpenStudy (asnaseer):

the denominator of your second term above is wrong

OpenStudy (asnaseer):

you should get:\[x^2/ x^2 + x^2/x = 6x^2\]

OpenStudy (asnaseer):

next - simplify each term

OpenStudy (anonymous):

but dont they need to have the same denominator to add

OpenStudy (asnaseer):

what does \(x^2/x^2\) equal?

OpenStudy (anonymous):

1

OpenStudy (asnaseer):

correct - now simplify the next term

OpenStudy (anonymous):

1+ x= 6x^2

OpenStudy (asnaseer):

thats right - now re-arrange this into a standard form, i.e into the form \(ax^2+bx+c=0\) and then solve

OpenStudy (anonymous):

so I did quadratic equation and go -1 plus or minus radical 25 all over -12

OpenStudy (asnaseer):

doesn't look right - can you please show your steps so that I can help spot where you may have made a mistake?

OpenStudy (anonymous):

k, u said put it into the standard form so I started from 1 + x = 6x^2 -6x^2 + x+1 = 0 a= -6, b= 1, c= 1 \[-1\pm \sqrt{1^{2}-4(-6)(1) \over 2(-6)}\]

OpenStudy (anonymous):

then got -1 plus or minus radical 25 all over -12

OpenStudy (asnaseer):

oh - sorry - I misread

OpenStudy (asnaseer):

you are right - and you know what \(\sqrt{25}\) equals?

OpenStudy (asnaseer):

the formula you used doesn't look quite right, it should be:\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (asnaseer):

the WHOLE thing should be divided by 2a

OpenStudy (asnaseer):

so you should have got:\[\frac{-1\pm \sqrt{1^{2}-4(-6)(1)}}{2(-6)}\]

OpenStudy (anonymous):

o ok so um x = 1/-3 and x = 1/2

OpenStudy (asnaseer):

perfect!

OpenStudy (anonymous):

ya I ment to put all over 2a but I was trying to make it come out nice and pretty in the equation thing but it took lng

OpenStudy (asnaseer):

although you should write it as: x=-1/3 or 1/2

OpenStudy (asnaseer):

you don't usually put the minus sign in the denominator

OpenStudy (anonymous):

wow these are so frustrating I was literally about to give up and go home, im in a library trying to study and there are all these other students basically hanging out and talking and there are no available tutors

OpenStudy (asnaseer):

also note that you should use the word "or" rather than "and" here. i.e. x = -1/3 or 1/2 instead of x = -1/3 and 1/2 because x canot equal both values at the same time

OpenStudy (anonymous):

ok ill do that

OpenStudy (anonymous):

thanks a lot you gave me a little hope to keep pushing a little more

OpenStudy (asnaseer):

great - and keep persevering - you'll get the hang of these very soon I'm sure :)

OpenStudy (asnaseer):

yw :)